# Spectroscopy using a visible photonic lantern at the Subaru telescope: Laboratory characterization and first on-sky demonstration on Ikiiki ( Leo) and ‘Aua ( Ori)

### v1

@@ 2.2.2 The spectrograph
The spectrograph is shown in Figure 4 (Right). The 19 SM fibers of the PL output are spliced with SM600 type SMFs. The SMFs (hereafter referred to as PL outputs) are rearranged in a linear V-groove (VG) which aligns the fibers side by side in a pseudo-slit feeding the spectrograph. The distance between consecutive PL outputs is µm for a total length of mm.
A 2 apochromatic microscope objective (Obj) with a 0.1 numerical aperture collimates the beams coming from the PL outputs. The collimated beams are then dispersed using a volume phase holographic (VPH) grating (manufactured by Wasatch Photonics). The transmission consists of two AR-coated glass covers protecting a layer of dichromated gelatin into which a lines.mm-1 interference pattern is written using coherent laser light. The VPH exhibits low scatter and low wavefront distortion. The manufacturer specifies that its dispersion efficiency, characterized between and nm, is polarization-independent and ranges from % to %. The grating is followed by a system of two imaging achromatic doublets of focal lengths of mm (AD1) and mm (AD2).
Spectra are imaged on a high frame rate, low readout noise camera (Hamamatsu ORCA-Quest qCMOS). It contains square pixels of m, a gain of about 0.11 /ADU, and its quantum efficiency ranges from to between and nm. Finally, a Wollaston prism (W) is placed upstream of the grating to separate the spectra associated with two orthogonal linear polarizations on the camera.
The resulting image is composed of 38 spectra coming from both polarized beams of the 19 SM outputs of the PL, as shown in Fig. 3. The spectra can then be combined (co-added) to reconstruct the source bulk spectrum.
@@ 3.1 Photonic lantern properties
The images and transmission measurements shown in Figure 5 were recorded using a laser diode tuned at 780nm (Thorlabs LPS-785-FC) with a measured standard deviation power stability of (over 2 hours) during the measurements. The transmissions values recorded in Figure 5 were obtained by using a true cut-back method commonly used in optical fiber components. To perform the cut-back loss measurements the output fiber of the laser was spliced to a SM PL pigtail and the output power at the PL MM end was recorded in a large area photo detector (Thorlabs S120C) by direct butt-coupling. Once the power was recorded, the SM PL pigtail was cleaved after the initial laser-SM PL splice and the power at the laser fiber end was recorded. The value of the transmission then is direct representation of the true transmission losses of the whole device by comparing those two measurements directly. This method is conveniently unaffected by input coupling efficiency from laser to the SM PL, nor on the splice losses of the measurement assembly. It is important to note that the presented values are raw values of this method and do not include the possible Fresnel reflections at the MM PL end during the measurements – this could be as high as on the silica-air interface. The recorded laser stability during the measurements and the significant figures of the large area photo detector apparatus gave us an error of on the recorded transmissions.
@@
The average measured transmission is , with values spanning between and . This is in accordance with the expected transmission expected in a quasi-adiabatic photonic lantern transition between MM and SM ends (Birks et al., 2015).
@@ 3.2.1 Injection optimization using coupling maps
The injection optimization into the PL is performed by scanning the focal plane in the (X,Y) directions by using the translation stages mentioned in Section 2.2.1. A scene window is defined in units of translation stage steps with two parameters: total width of the scanned window and its sampling. The stages then position the PL at each sampled (X,Y) coordinate, and for each location, an image of the PL outputs is acquired with the spectrograph. After the scan is complete, coupling maps are computed, corresponding to the total flux acquired for each (X,Y) position. The source used for this test is the SuperK.
@@
This process is performed simultaneously for each output and each wavelength, providing coupling maps with the number of outputs (here, 19) and the number of spectral channels (here, 1896 spanning from nm to nm). In addition, we compile the total coupling map, defined as the sum of all 19 output maps, capturing the global behavior of the PL. Figure 6 presents the resulting coupling maps for the 19 outputs, at a single wavelength (765.5 nm). This wavelength was chosen because it presented the best SNR. The injection setup was fixed with a focal ratio of 4.
@@
All coupling maps exhibit high order structures, similarly to the light distribution imaged at the PL input when retro-injecting in each SM output (see Fig. 5). The summed map is smoother, but still granulated. The darkest ”spot” of the total map is fainter than the brightest ”spot”. The size of the map envelope is about 26 , which is comparable to the beam size of the fundamental mode (about ) convolved with the PSF (about at 765 nm with a focal ratio of 4).
@@
Once the coupling map is acquired, a 2D Gaussian function is fitted to the coupling map to find the coordinates of its center, corresponding to the optical axis, or FoV center. We then drive the PL stages to these coordinates. Alternatively, one can choose to position the PL at another location, for example to maximize the flux injected into the PL by choosing the coordinates corresponding to the brightest spot of the summed coupling map. The summed coupling map is granulated, which means that the throughput is not homogeneous across the field. Figure 10 presents the PL coupling maps summed over the 19 outputs, for 6 different wavelengths. The comparison of these maps shows that the granulation pattern changes with wavelength.
@@
Because the size of the PSF is small compared to the fundamental MM PL input mode (by a factor of about 7.5 at 765 nm), the number of modes of the input wavefront is too large compared to the output 19 modes. An over-sampled input mode field at the input of a few-mode fiber (i.e. much smaller input PSFs diameters than the guided mode field diameters of the few-mode fiber) will result into a spatially and wavelength coupling efficiency mismatch and dependence on the input coupling conditions. Thus leading to the overall granulation observed in the summed coupling map. Figure 8 shows the total coupling map granulation for different focal ratios. At larger the focal ratio, the coupling map is less granulated. This is explained by the fact that a larger PSF (or larger focal ratio) corresponds to fewer modes sampled at the PL injection.
@@ 3.2.2 Field of View
The coupling maps are not only a proxy to optimize the injection into the PL, they can also inform on the FoV when converting the measured coupling maps from linear motion coordinates to sky angular coordinates. We subsequently compute the radial average of the resulting map to estimate angular field of view. The resulting curves, shown in Figure 10, have been normalized and show that the FoV increases when the focal ratio decreases. This can be taken into account when choosing the science application for the PL.
@@ 3.2.3 Parametric analysis of injected modes via singular value decomposition
We apply a singular value decomposition (SVD) to the coupling maps provided by the 19 outputs of the PL, for each tested focal ratio. To do so, each of the 19 coupling maps of dimension [,] is converted into a vector of dimension . We then construct a matrix of dimension []. The SVD is applied on the resulting matrix and provides the SVD modes (that we name injection modes) and singular values associated to each mode.
The modes are shown in Figure 9, noted from 1 to 19, for focal ratios ranging from 4 to 28, and for one wavelength ( = 765.5 nm). Apart from the first mode, which is an average of the global envelope, we see that the modes have more high frequency structures when the focal ratio decreases (hence, the PSF size decreases). For focal ratio values between 4 and 11.7, the lantern couples 19 input modes. For a focal ratio value of 28, there are only roughly 7 modes injected, and modes 8 to 19 seem to be mostly noise - consistent with the low singular values. Similar behavior is observed for the rest of the spectrum.
@@
We can also deduce from the graph that the matrix conditioning number — defined as the ratio of the largest singular value to the smallest singular value — increases with the focal ratio. This indicates that the stability of the inversion diminishes as we progressively introduce fewer ”genuine” input modes (or as the f-ratio increases). Among these modes at high f-ratio, some are authentic modes observed by the PL, while others merely constitute noise, as discussed below.
This modal study is a stepping stone towards a linear representation of the PL, and will be used in the future to investigate image reconstruction using the PL.
@@ 3.3.1 Injection versus wavelength
We now study the spectrum reconstruction of data acquired with the PL on the SCExAO super-continuum calibration source. We followed the previously mentionned injection optimization and placed the PL input at the center of the coupling map for each case. Depending on the application and noise regime, different spectrum extraction methods can be considered: from simply co-adding all the traces to performing a weighted signal extraction on each output, where only the bright parts of each trace are summed. In the photon-noise limited regime a simple co-addition of the traces is optimal. In a read-out-noise limited regime, minimizing the number of pixels leads to higher signal-to-noise (SNR), so a weighted signal extraction method should be considered. For the purpose of this paper, we will not discuss the various spectral extraction methods and simply choose to extract and co-add the 38 traces. This is especially suitable for objects characterization through spectroscopy, but also for Radial Velocity measurements. We also acquired a reference spectrum using the SMF mounted on the same injection setup as the PL. We present the reconstructed data in Figure 12.
Comparing the PL reconstructed spectra to the reference spectrum reveals some low frequency modulation for focal ratios 4 and 6. These correspond to a loss of flux between the input beam and the output of the PL. In the case of low focal ratios, due to the small PSF size there is more energy in high frequency spatial modes that are not supported by the 19 modes coupled from the MM input to the SM outputs, and hence cannot be injected leading to wavelength dependence losses in the overall spectra. This effect needs to be taken into account since it could make the continuum reconstruction challenging depending on the targeted science case. The spectra acquired for larger focal ratio do not exhibit these large modulation patterns: the spectral reconstruction is more accurate.
Some high frequency fringes are also observed in all of the PL spectra, but are not present in the SMF reference spectrum. This interference pattern, also visible in the raw data (not co-added spectra of the SMF outputs), is suspected to result from internal reflections on the PL and V-groove end faces (due to lack of an AR coating) having a Fabry-Perot effect, but requires further investigation.
Finally, the spectra reconstructed from the PL data show a decrease in intensity for wavelengths larger than about 760 nm, plus a hump at about 775 nm. This is due to the 800 nm short pass filter used at an angle for the pyramid WFS pickoff. The angle of the filter relative to the incident beam influences the transmission of both parallel and perpendicular polarizations of light in varying ways. The result is a bump located here around 770 nm in the spectral data, also visible in the SMF spectrum.
@@ 3.3.2 Injection efficiency versus focal ratio
Our procedure was similar for each tested focal ratio: we optimized the injection by acquiring a coupling map with the SCExAO SuperK and positioning the PL at the estimated center of the coupling map. Because our SuperK is not very bright in the visible wavelength, we needed to use a more powerful visible laser in order to make flux measurements using a power meter placed in the beam before and after the PL. After optimization on the SuperK, we swapped the source for a 642 nm laser, and recorded the flux intensity, with the power meter, before () and after () the PL. We normalized out the Fresnel reflections () at the input and output of the PL ( per surface), plus the average throughput () of the PL () in order to compute the coupling injection efficiency () following : . Owing to the granulation in the coupling maps, the coupling efficiency might not be the same over the whole FoV. Using the on-axis measurement, we converted the coupling map computed at nm into a coupling efficiency map by extrapolating the on-axis value to the rest of the map. From this coupling efficiency map, we compute two more quantities: the maximum coupling efficiency and the averaged injection efficiency over the FoV. We choose to define the FoV as the angular separation for which the injected flux in the lantern is half of the normalized flux in Figure 10.
Figure 13 summarizes our tests, highlighting the on-axis, maximum and averaged injection efficiency for various focal ratios. The error bars of the averaged injection efficiency represent the standard deviation of the coupling efficiency over the FoV. The injection efficiency shows a relatively fast increase from focal ratios 4 to 8. The highest injection efficiency is obtained for the focal ratio of 8, with an average value of over a field of view of about 80 mas, an on-axis injection efficiency of , and a maximum injection efficiency of . The injection efficiency then decreases for focal ratios ranging from 8 to 28. This trend was observed in Lin et al. (2022) when studying the coupling efficiency into a 19-port PL optimized for near-IR wavelengths. The graph also shows the FoV evolution as a function of the focal ratio. The curve follows a hyperbola. This is expected since the acceptance angle of a fiber is proportional to the inverse of the focal ratio.
@@ 3.3.3 Injection efficiency versus Strehl ratio
To assess the injection efficiency into the PL in the presence of uncorrected atmospheric turbulence, turbulence is injected onto the SCExAO DM with various levels of upstream atmospheric turbulence correction. Turbulence screens are based on Kolmogorov spectrum with inner and outer scales, adopting the frozen flow approximation for temporal evolution, and are the closest to what we expect during on-sky observations. We fixed the wind speed at 10 m/s and modified the turbulence amplitude to vary the Strehl ratio. In order to study the behavior of the PL in various conditions, we varied the spatial frequency content of the simulated turbulence. We identified three distinct cases by adjusting the inner and outer scales of the turbulence. First, turbulence following the Kolmogorov power spectrum (inner scale = 0.01 meter, outer scale = 20 meters), which is the closest to what we would commonly experience on SCExAO. Second, turbulence dominated by high spatial frequencies (inner scale = 0.01 meter, outer scale = 1 meter), which would correspond to an ExAO case where low order aberrations are well corrected but there is still remaining high order aberrations (for example if the DM lacks actuators). Third, turbulence dominated by the low spatial frequencies (inner scale = 10 meters, outer scale = 100 meters), which would correspond to a more classic AO loop performing poorly on low order aberrations (for example, during the observation of a faint star).
@@
We chose to conduct this study with the injection set at a focal ratio of about 8, where our on-axis injection efficiency was at 642 nm. The result is presented in Figure 14. The Strehl ratio was estimated from focal plane images simultaneously recorded on the VAMPIRES instrument at 750 nm. Our reference point was defined as the condition where there is no simulated turbulence on the bench.
@@
The estimated Strehl ratio without simulated turbulence was estimated at from the VAMPIRES images, which we associated with the injection efficiency measured on the bench. For each tested Strehl ratio, a total of 1000 PL frames were acquired. We performed the spectrum reconstruction from each frame following the same process than described in Section 3.3.1. We computed the average and standard deviation of the flux over the 1000 frames for each spectral channel. We then compared the averaged values for the 642 nm spectral channel to our reference to estimate the injection efficiency. Additionally, we mapped the evolution of the injection efficiency as a function of Strehl ratio for a simulated SMF. This simulation was conducted by Lin et al. (2021) in the context of designing PLs for ground-based telescopes and, among other aspects, involved comparing their theoretical injection efficiency with that of an SMF. If the simulations were run based on synthetic images of the Keck telescope, the results could easily be scaled to the Subaru telescope. A study by Jovanovic et al. (2017) demonstrated that the injection efficiency into an SMF on the Subaru telescope should be of that on the Keck telescope, due to the change of pupil shape. The graph also displays the various turbulence regimes (seeing, AO and ExAO).
@@
The graph indicates that the injection efficiency into the PL exhibits a linear relationship with Strehl ratio across all turbulence regimes in the case where high frequency contents are favored in the turbulence simulation. This behavior is similar to the one of the SMF, and can be explained by the non-sensitivity of the PL (or the SMF) to features at large separations. When the turbulence is dominated by low frequency aberrations, the injection to Strehl relationship remains linear but exhibits a slower decrease in efficiency with the Strehl degradation compared to the high-frequency dominated case. In the regime where the turbulence follows a classic Kolmogorov spectrum, the linearity is similar to that observed with low-frequency trends for both AO and ExAO turbulence regimes, until there is a drop in the seeing regime. This latter decline is attributed to tip/tilt dominating aberrations, as explained and simulated in Lin et al. (2021).
@@
This study reveals that the PL behaves similarly to a SMF, in terms of injection efficiency, when high order aberrations dominate. In a turbulence regime dominated by lower spatial frequency aberrations, the injection efficiency decreases slower with Strehl ratio degradation. In this latter regime, a PL provides higher coupling efficiency than an SMF, especially for Strehl ratios below 0.8. In practice, a PL and an SMF will have similar behaviors in the ExAO regime - with an advantage given to the PL especially in the presence of small residual tip/tilt. In the AO regime, a PL will have a better injection efficiency compared to an SMF.

### v3

@@ 2.2.2 The spectrograph
The spectrograph is shown in Fig. 4 (Right). The 19 SM fibers of the PL output are spliced with SM600 type SMFs. The SMFs (hereafter referred to as PL outputs) are rearranged in a linear V-groove (VG) that aligns the fibers side by side in a pseudo-slit feeding the spectrograph. The distance between consecutive PL outputs is µm for a total length of mm.
A 2 apochromatic microscope objective (Obj) with a 0.1 numerical aperture collimates the beams coming from the PL outputs. The collimated beams are then dispersed using a volume phase holographic (VPH) grating (manufactured by Wasatch Photonics). The transmission consists of two AR-coated glass covers protecting a layer of dichromated gelatin into which a lines.mm-1 interference pattern is written using coherent laser light. The VPH exhibits low scatter and low wavefront distortion. The manufacturer specifies that its dispersion efficiency, characterized between and nm, is polarization-independent and ranges from % to %. The grating is followed by a system of two imaging achromatic doublets with focal lengths of mm (AD1) and mm (AD2).
Spectra are imaged on a high-frame-rate, low-readout-noise camera (Hamamatsu ORCA-Quest qCMOS). It contains square pixels of m, a gain of about 0.11 /ADU, and its quantum efficiency ranges from to between and nm. Finally, a Wollaston prism (W) is placed upstream of the grating to separate the spectra associated with two orthogonal linear polarizations on the camera.
The resulting image is composed of 38 spectra coming from both polarized beams of the 19 SM outputs of the PL, as is shown in Fig. 3. The spectra can then be combined (co-added) to reconstruct the source bulk spectrum.
@@ 3.1 Photonic lantern properties
The images and transmission measurements shown in Fig. 5 were recorded using a laser diode tuned at 780nm (Thorlabs LPS-785-FC) with a measured standard deviation power stability of (over 2 hours) during the measurements. The transmissions values recorded in Fig. 5 were obtained by using a true cut-back method commonly used in optical fiber components. To perform the cut-back loss measurements, the output fiber of the laser was spliced to an SM PL pigtail and the output power at the PL MM end was recorded in a large area photo detector (Thorlabs S120C) by direct butt-coupling. Once the power was recorded, the SM PL pigtail was cleaved after the initial laser-SM PL splice and the power at the laser fiber end was recorded. The value of the transmission then is direct representation of the true transmission losses of the whole device by comparing those two measurements directly. This method is conveniently unaffected by either input coupling efficiency from the laser to the SM PL, or by the splice losses of the measurement assembly. It is important to note that the presented values are raw values of this method and do not include the possible Fresnel reflections at the MM PL end during the measurements – this could be as high as on the silica-air interface. The recorded laser stability during the measurements and the significant figures of the large area photo detector apparatus gave us an error of on the recorded transmissions.
@@
The average measured transmission is , with values spanning between and . This is in accordance with the expected transmission expected in a quasi-adiabatic PL transition between MM and SM ends (Birks et al., 2015).
@@ 3.2.1 Injection optimization using coupling maps
The injection optimization into the PL was performed by scanning the focal plane in the (X,Y) directions by using the translation stages mentioned in Sect. 2.2.1. A scene window was defined in units of translation stage steps with two parameters: the total width of the scanned window and its sampling. The stages then positioned the PL at each sampled (X,Y) coordinate, and for each location, an image of the PL outputs was acquired with the spectrograph. After the scan was complete, coupling maps were computed, corresponding to the total flux acquired for each (X,Y) position. The source used for this test was the SuperK.
@@
This process was performed simultaneously for each output and each wavelength, providing coupling maps, with the number of outputs (here, 19) and the number of spectral channels (here, 1896 spanning from nm to nm). In addition, we compiled the total coupling map, defined as the sum of all 19 output maps, capturing the global behavior of the PL. Figure 6 presents the resulting coupling maps for the 19 outputs, at a single wavelength (765.5 nm). This wavelength was chosen because it presented the best S/N. The injection setup was fixed with a focal ratio of 4.
@@
All coupling maps exhibit high-order structures, similarly to the light distribution imaged at the PL input when retro-injecting in each SM output (see Fig. 5). The summed map is smoother, but still granulated. The darkest “spot” of the total map is fainter than the brightest spot. The size of the map envelope is about 26 , which is comparable to the beam size of the fundamental mode (about ) convolved with the PSF (about at 765 nm with a focal ratio of 4).
@@
Once the coupling map was acquired, a 2D Gaussian function was fit to the coupling map to find the coordinates of its center, corresponding to the optical axis, or the FoV center. We then drove the PL stages to these coordinates. Alternatively, one can choose to position the PL at another location; for example, to maximize the flux injected into the PL by choosing the coordinates corresponding to the brightest spot of the summed coupling map. The summed coupling map is granulated, which means that the throughput is not homogeneous across the field. Figure 7 presents the PL coupling maps summed over the 19 outputs, for six different wavelengths. The comparison of these maps shows that the granulation pattern changes with wavelength.
@@
Because the size of the PSF is small compared to the fundamental MM PL input mode (by a factor of about 7.5 at 765 nm), the number of modes of the input wavefront is too large compared to the output 19 modes. An over-sampled input mode field at the input of a few-mode fiber (i.e., much smaller input PSFs diameters than the guided mode field diameters of the few-mode fiber) will result in a spatially and wavelength coupling efficiency mismatch and dependence on the input coupling conditions, leading to the overall granulation observed in the summed coupling map. Figure 8 shows the total coupling map granulation for different focal ratios. The larger the focal ratio, the less granulated the coupling map. This is explained by the fact that a larger PSF (or larger focal ratio) corresponds to fewer modes sampled at the PL injection.
@@ 3.2.2 Field of view
The coupling maps are not only a proxy to optimize the injection into the PL; they can also inform on the FoV when converting the measured coupling maps from linear motion coordinates to sky angular coordinates. We subsequently computed the radial average of the resulting map to estimate the angular FoV. The resulting curves, shown in Fig. 9, have been normalized and show that the FoV increases when the focal ratio decreases. This can be taken into account when choosing the science application for the PL.
@@ 3.2.3 Parametric analysis of injected modes via singular value decomposition
We applied a singular value decomposition (SVD) to the coupling maps provided by the 19 outputs of the PL, for each tested focal ratio. To do so, each of the 19 coupling maps of dimension [,] was converted into a vector of dimension . We then constructed a matrix, , with dimensions []. The SVD was applied on the resulting matrix, , and provides the SVD modes (that we name injection modes) and singular values associated with each mode.
The modes are shown in Fig. 10, noted from 1 to 19, for focal ratios ranging from 4 to 28, and for one wavelength ( = 765.5 nm). Apart from the first mode, which is an average of the global envelope, we see that the modes have more high-frequency structures when the focal ratio decreases (hence, the PSF size decreases). For focal ratio values between 4 and 11.7, the lantern couples 19 input modes. For a focal ratio value of 28, there are only roughly seven modes injected, and modes 8 to 19 seem to be mostly noise – consistent with the low singular values. Similar behavior is observed for the rest of the spectrum.
@@
We can also deduce from the graph that the matrix conditioning number — defined as the ratio of the largest singular value to the smallest singular value — increases with the focal ratio. This indicates that the stability of the inversion diminishes as we progressively introduce fewer “genuine” input modes (or as the f-ratio increases). Among these modes at high f-ratio, some are authentic modes observed by the PL, while others merely constitute noise, as is discussed below. This modal study is a stepping stone toward a linear representation of the PL, and will be used in the future to investigate image reconstruction using the PL.
@@ 3.3.1 Injection versus wavelength
We now study the spectrum reconstruction of data acquired with the PL on the SCExAO super-continuum calibration source. We followed the previously mentioned injection optimization and placed the PL input at the center of the coupling map in each case. Depending on the application and noise regime, different spectrum extraction methods can be considered: from simply co-adding all the traces to performing a weighted signal extraction on each output, where only the bright parts of each trace are summed. In the photon-noise limited regime, a simple co-addition of the traces is optimal. In a read-out-noise-limited regime, minimizing the number of pixels leads to higher signal-to-noise (S/N), so a weighted signal extraction method should be considered. For the purpose of this paper, we shall not discuss the various spectral extraction methods, and have simply chosen to extract and co-add the 38 traces. This is especially suitable for objects characterization through spectroscopy, but also for radial velocity measurements. We also acquired a reference spectrum using the SMF mounted on the same injection setup as the PL. We present the reconstructed data in Fig. 12.
Comparing the PL reconstructed spectra to the reference spectrum reveals some low-frequency modulation for focal ratios 4 and 6. These correspond to a loss of flux between the input beam and the output of the PL. In the case of low focal ratios, due to the small PSF size there is more energy in high-frequency spatial modes that are not supported by the 19 modes coupled from the MM input to the SM outputs, and hence cannot be injected leading to wavelength dependence losses in the overall spectra. This effect needs to be taken into account since it could make the continuum reconstruction challenging depending on the targeted science case. The spectra acquired for larger focal ratio do not exhibit these large modulation patterns: the spectral reconstruction is more accurate.
Some high-frequency fringes are also observed in all of the PL spectra, but are not present in the SMF reference spectrum. This interference pattern, also visible in the raw data (not co-added spectra of the SMF outputs), is suspected to result from internal reflections on the PL and VG end faces (due to the lack of an AR coating) having a Fabry-Perot effect, but requires further investigation.
Finally, the spectra reconstructed from the PL data show a decrease in intensity for wavelengths larger than about 760 nm, plus a hump at about 775 nm. This is due to the 800 nm short pass filter used at an angle for the pyramid WFS pickoff. The angle of the filter relative to the incident beam influences the transmission of both parallel and perpendicular polarizations of light in varying ways. The result is a bump located here around 770 nm in the spectral data, which is also visible in the SMF spectrum.
@@ 3.3.2 Injection efficiency versus focal ratio
Our procedure was similar for each tested focal ratio: we optimized the injection by acquiring a coupling map with the SCExAO SuperK and positioning the PL at the estimated center of the coupling map. Because our SuperK is not very bright in the visible wavelength, we needed to use a more powerful visible laser in order to make flux measurements using a power meter placed in the beam before and after the PL. After optimization on the SuperK, we swapped the source for a 642 nm laser, and recorded the flux intensity, with the power meter, before () and after () the PL. We normalized out the Fresnel reflections () at the input and output of the PL ( per surface), plus the average throughput () of the PL () in order to compute the coupling injection efficiency (), following . Owing to the granulation in the coupling maps, the coupling efficiency might not be the same over the whole FoV. Using the on-axis measurement, we converted the coupling map computed at nm into a coupling efficiency map by extrapolating the on-axis value to the rest of the map. From this coupling efficiency map, we computed two more quantities: the maximum coupling efficiency and the averaged injection efficiency over the FoV. We chose to define the FoV as the angular separation for which the injected flux in the lantern is half of the normalized flux in Fig. 7.
Figure 13 summarizes our tests, highlighting the on-axis, maximum, and averaged injection efficiency for various focal ratios. The error bars of the averaged injection efficiency represent the standard deviation of the coupling efficiency over the FoV. The injection efficiency shows a relatively fast increase from focal ratios 4 to 8. The highest injection efficiency is obtained for the focal ratio of 8, with an average value of over an FoV of about 80 mas, an on-axis injection efficiency of , and a maximum injection efficiency of . The injection efficiency then decreases for focal ratios ranging from 8 to 28. This trend was observed in Lin et al. (2022) when studying the coupling efficiency into a 19-port PL optimized for near-infrared wavelengths. The graph also shows the FoV evolution as a function of the focal ratio. The curve follows a hyperbola. This is expected since the acceptance angle of a fiber is proportional to the inverse of the focal ratio.
@@ 3.3.3 Injection efficiency versus Strehl ratio
To assess the injection efficiency into the PL in the presence of uncorrected atmospheric turbulence, turbulence is injected onto the SCExAO DM with various levels of upstream atmospheric turbulence correction. Turbulence screens are based on Kolmogorov spectrum with inner and outer scales, adopting the frozen flow approximation for temporal evolution, and are the closest to what we expect during on-sky observations. We fixed the wind speed at 10 m/s and modified the turbulence amplitude to vary the Strehl ratio. In order to study the behavior of the PL in various conditions, we varied the spatial frequency content of the simulated turbulence. We identified three distinct cases by adjusting the inner and outer scales of the turbulence. First, turbulence following the Kolmogorov power spectrum (inner scale = 0.01 meter, outer scale = 20 meters), which is the closest to what we would commonly experience on SCExAO. Second, turbulence dominated by high spatial frequencies (inner scale = 0.01 meter, outer scale = 1 meter), which would correspond to an ExAO case where low-order aberrations are well corrected but there are still remaining high-order aberrations (for example, if the DM lacks actuators). Third, turbulence dominated by the low spatial frequencies (inner scale = 10 meters, outer scale = 100 meters), which would correspond to a more classic AO loop performing poorly on low-order aberrations (for example, during the observation of a faint star).
@@
We chose to conduct this study with the injection set at a focal ratio of about 8, where our on-axis injection efficiency was at 642 nm. The result is presented in Fig. 14. The Strehl ratio was estimated from focal plane images simultaneously recorded on the VAMPIRES instrument at 750 nm. Our reference point was defined as the condition where there is no simulated turbulence on the bench.
@@
The estimated Strehl ratio without simulated turbulence was estimated at from the VAMPIRES images, which we associated with the injection efficiency measured on the bench. For each tested Strehl ratio, a total of 1000 PL frames were acquired. We performed the spectrum reconstruction from each frame following the same process than described in Sect. 3.3.1. We computed the average and standard deviation of the flux over the 1000 frames for each spectral channel. We then compared the averaged values for the 642 nm spectral channel to our reference to estimate the injection efficiency. Additionally, we mapped the evolution of the injection efficiency as a function of Strehl ratio for a simulated SMF. This simulation was conducted by Lin et al. (2021) in the context of designing PLs for ground-based telescopes and, among other aspects, involved comparing their theoretical injection efficiency with that of an SMF. If the simulations were run based on synthetic images of the Keck telescope, the results could easily be scaled to the Subaru Telescope. A study by Jovanovic et al. (2017) demonstrated that the injection efficiency into an SMF on the Subaru Telescope should be of that on the Keck telescope, due to the change in the pupil shape. The graph also displays the various turbulence regimes (seeing, AO, and ExAO).
@@
The graph indicates that the injection efficiency into the PL exhibits a linear relationship with the Strehl ratio across all turbulence regimes in the case in which high-frequency contents are favored in the turbulence simulation. This behavior is similar to the one of the SMF, and can be explained by the non-sensitivity of the PL (or the SMF) to features at large separations. When the turbulence is dominated by low-frequency aberrations, the injection to Strehl relationship remains linear but exhibits a slower decrease in efficiency with the Strehl degradation compared to the high-frequency dominated case. In the regime in which the turbulence follows a classic Kolmogorov spectrum, the linearity is similar to that observed with low-frequency trends for both AO and ExAO turbulence regimes, until there is a drop in the seeing regime. This latter decline is attributed to tip-tilt dominating aberrations, as is explained and simulated in Lin et al. (2021).
@@
This study reveals that the PL behaves similarly to a SMF, in terms of injection efficiency, when high-order aberrations dominate. In a turbulence regime dominated by lower-spatial-frequency aberrations, the injection efficiency decreases slower with Strehl ratio degradation. In this latter regime, a PL provides higher coupling efficiency than an SMF, especially for Strehl ratios below 0.8. In practice, a PL and an SMF will have similar behaviors in the ExAO regime - with an advantage given to the PL especially in the presence of small residual tip or tilt. In the AO regime, a PL will have a better injection efficiency compared to an SMF.